OpenStudy (anonymous):

3 divided by x exponent 2 + 2x + 1 subtract 9 divided by (x-3)

OpenStudy (studygurl14):

Like this? \(\large\frac{3}{x^2}+2+2x+1-\frac{9}{x-3}\)

OpenStudy (anonymous):

what do you think it is SilverPersia

OpenStudy (studygurl14):

@SilverPersia you there?

OpenStudy (anonymous):

yeah thats the problem @StudyGurl14

OpenStudy (leonardo0430):

whats the question :l

OpenStudy (studygurl14):

Okay, so first off combine the like terms (in blue). \(\large\frac{3}{x^2}\color{blue}{+2}+2x\color{blue}{+1}-\frac{9}{x-3}\)